OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 12, c = 19 14.03 14.38 14.73 15.07

OpenStudy (anonymous):

c

jimthompson5910 (jim_thompson5910):

Use the pythagorean theorem to get a^2 + b^2 = c^2 a^2 + 12^2 = 19^2 a^2 + 144 = 361 ... ... ... a = ???

OpenStudy (anonymous):

@Nylyak got what @jim_thompson5910 said right?

OpenStudy (anonymous):

\[a=\sqrt{361-144}\] which becomes: \[a=\sqrt{217}\]